OpenStudy (anonymous):

rs = 8y +4, st = 4y + 8 and rt = 15y - 9

OpenStudy (anonymous):

You'd have to find out what Y is first. This is like .. transitive property or something, so .. basically, RS and ST are going to equal RT. Soo .. 8y + 4 + 4y + 8 = 15y - 9. Simplify. 12y +12 = 15y - 9 Add 9 on both sides. 12y + 21 = 15y Subtract 12y from both sides. 21 = 3y Divide 3 on both sides. 7 = y y = 7 So ... Back to our original stuff. RS = 8y + 4 ST = 4y + 8 and RT = 15y - 9. Plug in "7" fr "y". RS = 8(7) + 4 ST = 4(7) + 8 and RT = 15(7) - 9 There you go.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yup